# Foundations of Reasoning in Language Models (FoRLM)

### v2

@@ 3.2.3 Memory Consolidation and Knowledge Evolution
4 Discussion
Our algorithmic formalisations of Flavell’s and Nelson and Narens’ frameworks offer one possible interpretation of reasoning system limitations. The prefix dominance trap might be understood as reflecting absent pre-generation monitoring – without FOK assessments and strategy selection, systems cannot evaluate approaches before committing. Hallucination could correspond to Nelson and Narens’ commission errors, suggesting missing confidence thresholds or termination criteria. Inconsistent self-evaluation may indicate the absence of hierarchical monitoring structures that maintain coherence in human metacognition. While these formalisations provide structured ways to think about computational components – monitoring loops, confidence transformations, working memory structures – we acknowledge that psychological theories developed for human cognition may not map cleanly onto artificial systems. The following specifications suggest architectural considerations rather than definitive solutions, and considerable work remains to determine whether these cognitive principles can meaningfully guide computational implementation.
Dynamic Memory-Guided Reasoning Against Hallucination Nelson and Narens’ $\lambda_{\text{confidence}}$ mechanism suggests a theoretical approach to hallucination through coordinated memory-reasoning confidence assessment. Their framework implies that systems might maintain domain-specific thresholds: $\lambda_{\text{confidence}}^{\text{(domain)}}=\text{median}(\{\text{confidence}_{\tau}:\text{correct outputs in domain}\})$, calibrated from performance history. When confidence falls below $\lambda_{\text{confidence}}^{(\tau)}$, Nelson and Narens’ model prescribes continuing search rather than outputting uncertain answers – directly addressing the commission error problem underlying hallucination. The dynamic adjustment mechanism $\beta_{\tau}=\exp(-\alpha\cdot\text{search burden})$ suggests how systems might balance accuracy against computational cost, maintaining high thresholds for factual domains while adapting to task demands. This distinction between commission errors (hallucination) and omission errors (refusing to answer) provides theoretical grounding for confidence-based output control.
Attention Coordination Through Metacognitive Monitoring The finding that LLMs operate within a lower-dimensional “metacognitive space” (Ji-An et al., 2025) resonates with multiple theoretical predictions. Flavell’s $\mathcal{ME}_{\text{difficulty}}^{\tau}$ formulation suggests hierarchical monitoring of cognitive states, while Nelson and Narens’ FOK mechanism implies rapid accessibility assessment operating faster than full memory retrieval – consistent with Reder’s [1987; 1988] findings that FOK judgments have shorter latency than actual recall. This convergence suggests that efficient metacognitive monitoring might naturally operate on compressed representations of cognitive state. The theoretical framework indicates how different reasoning processes might require different levels of monitoring granularity, potentially explaining why models report some computations accurately while failing on others (Lindsey et al., 2025). These insights, combined with findings that implicit confidence measures show greater metacognitive sensitivity than explicit ones (Xiong et al., 2023), suggest that attention allocation could benefit from monitoring internal representations rather than relying solely on verbalized assessments.
Strategic Reasoning Selection via Metacognitive Control MeCo’s dual-threshold policy (Li et al., 2025) parallels both Flavell’s strategy selection mechanism and Nelson and Narens’ dual-counter FOK hypothesis. Flavell’s formulation $\mathcal{CS}_{\tau}=\text{select}(s\in\mathcal{MK}_{\text{Strategy}}\mid\mathcal{ME}_{\text{difficulty}}^{\tau},\mathcal{MK}_{\tau},\mathcal{T},\mathcal{G})$ suggests how metacognitive experiences guide strategy choice. Nelson and Narens’ use of FOK magnitude $\text{FOK}_{\tau}||$ for search intensity decisions – intensive attention when $\text{FOK}_{\tau}||<\lambda_{\text{FOK}}^{(\tau)}$ versus standard or termination based on counter dominance – provides a principled mechanism for adaptive control. MeCo’s implementation of weak versus strong metacognitive signals, detected through linearly separable directions in representation space (Zou et al., 2023; Liu et al., 2023), mirrors this magnitude-based approach. This convergence indicates how reasoning strategy selection – whether chain-of-thought, direct retrieval, or tool invocation – could be guided by formal metacognitive principles rather than heuristic rules.
Memory Consolidation for Adaptive Learning Nelson and Narens’ experience consolidation mechanism ($\Phi_{\tau}^{\text{STM}}\xrightarrow{\psi}\text{LTM}$) pinpoints a fundamental gap in current LLM capabilities. Their formulation of experience tuples $\Phi_{\tau}=(\mathcal{ME}_{\tau},\text{Strategy}_{\tau},\text{Outcome}_{\tau})$ suggests how systems might learn from their own cognitive experiences, updating metacognitive knowledge about strategy effectiveness, monitoring accuracy, and resource requirements. This aligns with Didolkar et al. (2024)’s extraction of metacognitive knowledge through skill annotations, though current approaches capture only static snapshots rather than dynamic learning. The framework’s emphasis on pattern recognition across experience tuples – identifying which strategies succeed for which task types under which confidence levels – suggests that true adaptive behavior requires explicit mechanisms for metacognitive memory beyond current in-context learning. While current LLMs rely on fixed architectures with static learning mechanisms, concepts like the Darwin Gödel Machine (Zhang et al., 2025) point toward systems that could empirically evolve their own metacognitive strategies, transforming the consolidation process itself based on accumulated experience.
5 Limitations
While our formalisation provides theoretical foundations for understanding reasoning failures through metacognitive principles, several limitations constrain the immediate applicability of this work.
Implementation challenges We have not proposed how to implement metacognitive constructs such as $\mathcal{ME}_{\text{difficulty}}$ or $\text{FOK}^{+/-}$ in practice. Whether these subjective “feelings” should be realised as learned representations, designed features, or emerge from architectural properties remains unspecified. The translation from psychological experiences to computational mechanisms presents non-trivial challenges, particularly given the parallel nature of transformer architectures versus the sequential processing assumed in our algorithms.
Computational complexity The MGV framework introduces substantial overhead through iterative monitoring-generation-verification cycles, multiple metacognitive state updates, and experience consolidation mechanisms. We have not analysed the computational cost-benefit trade-offs, nor determined whether simpler metacognitive mechanisms might achieve comparable benefits with reduced complexity.
Theoretical foundations We largely accept Flavell’s and Nelson and Narens’ theories without critical examination of their limitations or controversies within cognitive science. Alternative metacognitive theories might provide different insights, and the suitability of these particular frameworks for computational systems remains unexamined.
6 Conclusion
This paper has presented a theoretical exploration of how metacognitive principles from cognitive science might address fundamental limitations in language model reasoning architectures. Through systematic formalisation of Flavell’s cognitive monitoring model and Nelson and Narens’ metamemory framework, we have translated psychological constructs into computational structures, revealing specific architectural gaps in current test-time reasoning approaches. The resulting Monitor-Generate-Verify framework suggests how explicit metacognitive mechanisms – pre-generation monitoring, dynamic confidence thresholds, and experience consolidation – could potentially address issues such as the prefix dominance trap.
However, translating these theoretical insights into practical architectures remains an open challenge. The gap between abstract metacognitive experiences and implementable computational mechanisms is substantial, and whether the additional complexity of MGV architectures would yield commensurate benefits requires empirical investigation. Future work must address how to operationalise subjective monitoring signals, implement working memory mechanisms compatible with parallel processing, and design efficient consolidation processes that enable genuine metacognitive learning.
Despite these limitations, we believe examining established cognitive theories provides valuable perspectives on architectural evolution. As language models increasingly tackle complex reasoning tasks, understanding the computational principles underlying human metacognition may guide development of systems that monitor their own cognitive states, adaptively select strategies, and learn from their reasoning experiences. Whether through direct implementation of MGV principles or through architectures inspired by metacognitive insights, bridging cognitive science and artificial intelligence remains a promising direction for advancing reasoning capabilities.
Appendix A Flavell’s Model of Cognitive Monitoring
Flavell [1979] describes metacognition as a dynamic control architecture comprising four interacting components: metacognitive knowledge, metacognitive experience, goals (or tasks), and actions (or strategies). Rather than viewing metacognition as merely stored knowledge about cognition, Flavell presents it as a dynamic control system. This system operates through continuous interactions between four elements: what agents know about their cognitive capabilities (metacognitive knowledge), what they feel about their current cognitive state (metacognitive experiences), what they aim to achieve (goals), and how they control their thinking (strategies). Central to Flavell’s model is the principle of reciprocal interaction amongst these components. Rather than operating as independent modules, they form an integrated system characterised by continuous bidirectional influences: metacognitive knowledge guides both strategy selection and the interpretation of ongoing cognitive experiences; these conscious experiences, in turn, update the knowledge base and prompt strategic adjustments; task-goals determine which aspects of metacognitive knowledge become most salient; and the outcomes of chosen actions provide feedback that shapes both immediate metacognitive experiences and longer-term understanding of effective cognitive approaches. This dynamic interplay positions metacognition as a self-regulating system capable of adaptive control over cognitive processes.
Initialisation Let $\mathcal{T}$ be a task and $\mathcal{G}$ be the associated goal. We establish the initial system state:
$\mathcal{S}_{0}=f(\mathcal{T},\mathcal{G})$
where $(\mathcal{T},\mathcal{G)}$ is self-imposed or externally-imposed.
While Flavell [1979] treats ‘goals’ and ‘tasks’ as equivalent, we maintain a computational distinction to enhance the model’s precision: $\mathcal{T}$ represents the specific cognitive enterprise, whilst $\mathcal{G}$ represents the desired outcome or success criteria. This separation enables clearer analysis of metacognitive processes, such as assessing the cognitive demands of $\mathcal{T}$ relative to $\mathcal{G}$, or specifying which approaches to employ for $\mathcal{T}$ to achieve $\mathcal{G}$. For instance, the same reasoning task ($\mathcal{T}$: logical problem-solving) might require different metacognitive assessments depending on whether the goal is speed ($\mathcal{G}_{1}$: quick approximation) or accuracy ($\mathcal{G}_{2}$: verified solution).
M-G-V (Information Processing) Cycle For monitoring cycles $\tau=0,1,\ldots,T$:
WHILE $\mathcal{S}_{\tau}=\text{ACTIVE}$:
  1. 1.
    MONITOR: Monitor cognitive status through retrieval of metacognitive knowledge and assessment of metacognitive experience.
    Knowledge activation operates differently across metacognitive cycles, with initial cycles relying solely on task-goal combinations while ongoing cycles incorporate emerging metacognitive experiences. At $t=0$, the system identifies potentially relevant metacognitive knowledge based exclusively on the task-goal pairing. In subsequent cycles ($\tau>0$), the knowledge base expands as metacognitive experiences ($\mathcal{ME}$) from the previous cycle $\tau-1$ triggering additional relevant knowledge.
    $\mathcal{MK}_{\tau}=\begin{cases}\text{retrieve}(\mathcal{MK},\mathcal{T},\mathcal{G})&\text{if }\tau=0\\ \mathcal{MK}_{\tau-1}\cup\text{retrieve}(\mathcal{MK},\mathcal{ME}_{\tau-1})&\text{if }\tau>0\end{cases}$
    According to Flavell [1979], metacognitive knowledge comprises three major categories:
    • Agent Variables ($\mathcal{MK}_{\text{Agent}}$): Knowledge about cognitive agents’ characteristics and capabilities that applies across different cognitive endeavours. These are fundamentally subjective beliefs about processing preferences, strengths, and limitations rather than objective assessments. For computational agents, these may represent learned self-models – representations of performance patterns, processing preferences, and comparative capabilities derived from experience across cognitive tasks.
    • Task Variables ($\mathcal{MK}_{\text{Task}}$): Knowledge about cognitive situation assessment, including: (1) information characteristics (e.g., familiarity, complexity, organisation), and (2) task demands and goals. This knowledge is evaluative – understanding what task characteristics mean for cognitive processes and goal achievement, not merely recognising the characteristics themselves.
    • Strategy Variables ($\mathcal{MK}_{\text{Strategy}}$): Knowledge concerning the effectiveness of cognitive strategies ($\mathcal{CS}$) and metacognitive strategies ($\mathcal{MS}$). Across different goals and task types, $\mathcal{CS}$ are cognitive operations that address problem-solving procedures such as applying domain-specific algorithms or step-by-step problem decomposition, whereas $\mathcal{MS}$ monitor and regulate such cognitive processes. For instance, chain-of-thought reasoning represents a $\mathcal{CS}$ for solving problems systematically, while deciding to employ chain-of-thought based on problem complexity assessment represents a $\mathcal{MS}$. Flavell [1979] explicitly incorporates both strategy types within this category, reflecting his theoretical position that strategy selection constitutes a fundamentally metacognitive process requiring knowledge about when, how, and why particular approaches prove effective under specific conditions.
    These categories function as an integrated system: task variables diagnose cognitive demands, strategy variables prescribe responses, and agent variables contextualise both within the agent’s capabilities [^1] The distinction between $\mathcal{MK}_{\text{Strategy}}$ and $\mathcal{MK}_{\text{Task}}$ emerges from their functional roles. $\mathcal{MK}_{\text{Task}}$ enables diagnosis by identifying what makes cognitive enterprises demanding and how task characteristics influence goal achievement probability, whereas $\mathcal{MK}_{\text{Strategy}}$ enables prescription by specifying which cognitive approaches to deploy given those diagnostic assessments. Task variables answer ’what challenges does $\mathcal{T}$ present relative to $\mathcal{G}$?’ and strategy variables answer ’which approaches for $\mathcal{T}$ will achieve $\mathcal{G}$?’.
    Flavell [1979] distinguishes between knowledge-based experiences, which ‘are best described as items of metacognitive knowledge that have entered consciousness’ (e.g., suddenly recalling a relevant strategy), and feeling-based experiences, which ‘clearly cannot be described that way’ (e.g., feeling confused).
    This dual nature of metacognitive experience – alternating between immediate phenomenological feelings and knowledge-based assessments – motivates our formal representation using the exclusive-or operator $\mathcal{\oplus}$. In this formulation, $feel(·)$ captures the pure subjective sensations of cognitive state, while $assess(·)$ represents evaluations informed by metacognitive knowledge. The operator $\mathcal{\oplus}$ thus reflects Flavell’s distinction between feeling-based experiences (phenomenological states that cannot be reduced to knowledge) and knowledge-based experiences (instances of metacognitive knowledge entering consciousness).
    Our exclusive-or formalisation captures the observation that these two modes typically alternate rather than blend, though we acknowledge that this binary representation constitutes a modelling simplification of potentially richer interactions. Accordingly, this binary characterisation suggests that, at any given moment, an agent experiences either raw cognitive feelings awaiting interpretation or automatic, knowledge-influenced assessments. The temporal alternation between these exclusive states gives rise to the evolving metacognitive experience that guides subsequent processing.
    At this stage, it is notable that Flavell [1979, p. 909] primarily associates metacognitive experience with the subjective sense of perceived difficulty. Such experiences may involve feelings of complexity, comprehension challenges, conceptual opacity, or the sense that material exceeds current capabilities.
    $\mathcal{ME}_{\text{difficulty}}^{\tau}=\begin{cases}\text{feel}(\mathcal{T})\oplus\text{assess}(\mathcal{T},\mathcal{MK}_{\tau})&\text{if }\tau=0\\ \text{feel}(\mathcal{T},\text{Outcomes}_{\tau-1})\oplus\text{assess}(\mathcal{T},\text{Outcomes}_{\tau-1},\mathcal{MK}_{\tau})&\text{if }\tau>0\end{cases}$
    Accordingly, $\mathcal{ME}_{\text{difficulty}}^{\tau}$ evolves through iterative cycle-dependent assessments, progressing from initial coarse-grained feelings to increasingly nuanced evaluations that identify specific challenge sources and their implications for strategy selection. These experiences could help identify specific sources of obstacles and serve to guide the agent’s attentional and regulatory focus.
  2. 2.
    GENERATE: Control cognitive activity through strategy selection and execution.
    Flavell [1979, p. 909] emphasises that this stage centres on the selection of cognitive strategies ($\mathcal{CS}_{\tau}$) through the integration of metacognitive experiences and knowledge. Metacognitive experiences of difficulty ($\mathcal{ME}_{\text{difficulty}}$), whether feeling-based or knowledge-based, function as computational signals that indicate cognitive status. However, these signals require interpretation through metacognitive knowledge to guide effective strategy selection. As Flavell [1979, p. 906] establishes, effective cognitive regulation emerges only when metacognitive experiences combine with metacognitive knowledge, transforming ambiguous feelings into actionable strategic decisions.
    The strategy selection process draws upon $\mathcal{MK}_{\text{Strategy}}$, which encompasses knowledge about both metacognitive strategies ($\mathcal{MS}$) and cognitive strategies ($\mathcal{CS}$). Although Flavell does not specify the exact selection mechanism, his examples suggest a two-phase pattern-matching process. In the first phase, $\mathcal{MK}_{\text{Strategy}}$ guides the interpretation of $\mathcal{ME}_{\text{difficulty}}^{\tau}$, transforming general difficulty signals into precise diagnostic patterns. For instance, metacognitive knowledge might specify “when experiencing content uncertainty, identify specific unknown terms” or “when procedurally confused, assess whether confusion stems from missing steps versus unclear sequence”. In the second phase, these refined difficulty patterns activate corresponding cognitive strategies from $\mathcal{MK}_{\text{Strategy}}$ – content uncertainty with identified terms triggers seeking definitions, whilst procedural confusion from missing steps activates searching for worked examples. Throughout this process, the selection mechanism integrates agent capabilities ($\mathcal{MK}_{\text{Agent}}$) and task characteristics ($\mathcal{MK}_{\text{Task}}$) to identify the most appropriate strategy for achieving $\mathcal{G}$ given $\mathcal{T}$.
    $\mathcal{CS}_{\tau}=\text{select}(s\in\mathcal{MK}_{\text{Strategy}}\mid\mathcal{ME}_{\text{difficulty}}^{\tau},\mathcal{MK}_{\tau},\mathcal{T},\mathcal{G})$
    The selected cognitive strategy is implemented to produce cognitive outcomes ($\mathcal{CO}_{\tau}$), generating a feedback that is rich in nature, as as it encompasses not only task progress information but also a new context for the next cycle’s monitoring and potential strategy adjustment.
    $\mathcal{CO}_{\tau}=\text{execute}(\mathcal{CS}_{\tau},\mathcal{T},\mathcal{G})$
  3. 3.
    VERIFY: Evaluate progress and determine continuation.
    Following strategy execution, Flavell [1979, p. 909] comments that the outcomes potentially ‘trigger additional metacognitive experiences about how the endeavour is faring’. These evaluative experiences ($\mathcal{ME}_{\text{evaluative}}^{\tau}$) are about performance rather than difficulty.
    $\mathcal{ME}_{\text{evaluative}}^{\tau}=\text{feel}(\mathcal{\mathcal{CO}_{\tau}})\oplus\text{assess}(\mathcal{CO}_{\tau},\mathcal{MK}_{\tau})$
    These experiences, again informed and guided by pertinent metacognitive knowledge, instigate the metacognitive strategy of surveying ‘all that [the agent has] learned to see if it fits together into a coherent whole, if it seems plausible and consistent with [the agent’s] prior knowledge and expectations, and if it provides an avenue to the goal’ [Flavell, 1979, p. 909].
    $\mathcal{MS}_{\tau}=\text{select}(s\in\mathcal{MK}_{\text{Strategy}}^{\text{meta}}\mid\mathcal{ME}_{\text{evaluative}}^{\tau},\mathcal{MK}_{\tau},\mathcal{CO}_{\tau},\mathcal{G})$
    where
[…]
    $\mathcal{ME}_{\text{evaluative}}^{\tau}$ signals the need for assessment. For example, “feeling uncertain about validity of the processed outcome” or “sensing incomplete understanding despite completion of process”. These evaluative experiences activate relevant metacognitive strategies from $\mathcal{MK}_{\text{Strategy}}$. For instance, uncertainty about validity triggers plausibility checking, while sensing incompleteness activates coherence assessment to identify gaps.
    $\mathcal{MS}_{\tau}$ represents the strategic choice to conduct comprehensive evaluation along four possible dimensions: coherence (“do the outcomes form a consistent understanding?”), plausibility (“are the results believable given prior knowledge?”), consistency (“do outcomes align with initial expectations?”), and goal-conduciveness (“do current results provide a pathway to goal achievement?”). The execution systematically evaluates $\mathcal{CO}_{\tau}$ against relevant knowledge:
    $\mathcal{MO}_{\tau}=\text{execute}(\mathcal{MS}_{\tau},\mathcal{CO}_{\tau},\mathcal{MK}_{\tau},\mathcal{G})$
    Flavell emphasises that metacognitive experiences can ‘add to’, ‘delete from’, or ‘revise’ the metacognitive knowledge base through Piagetian [Flavell, 1963] mechanisms. The agent observes relationships among goals, strategies, experiences, and outcomes across the complete monitoring cycle.
    Let $\Phi_{\tau}$ represents the complete experience tuple, where $\mathcal{ME}_{\tau}=(\mathcal{ME}_{\text{difficulty}}^{\tau},\mathcal{ME}_{\text{evaluative}}^{\tau})$, $\text{Strategy}_{\tau}=(\mathcal{CS}_{\tau},\mathcal{MS}_{\tau})$ and $\text{Outcome}_{\tau}=(\mathcal{CO}_{\tau},\mathcal{MO}_{\tau})$:
    $\Phi_{\tau}$ $=(\mathcal{ME}_{\tau},\text{Strategy}_{\tau},\text{Outcome}_{\tau})$ (1)
    $\mathcal{MK}$ $=\text{update}(\mathcal{MK},\Phi_{\tau})$ (2)
    Based on the comprehensive metacognitive evaluation, the system determines its next state:
    $\mathcal{S}_{\tau+1}=\begin{cases}\text{ACTIVE}&\text{if }\neg\text{goal\_achieved}(\mathcal{CO}_{\tau},\mathcal{G})\\ \text{TERMINATE}&\text{if }\text{goal\_achieved}(\mathcal{CO}_{\tau},\mathcal{G})\end{cases}$
Appendix B Nelson and Narens’ Model of Metamemory
Nelson and Narens [1990] theorise metacognitive systems with particular focus on metamemory in the context of self-directed, self-paced learning and retrieval tasks. Their framework establishes metacognition as fundamentally hierarchical, distinguishing between cognitive processes that operate on mental content (object-level) and those that operate on cognitive processes themselves (meta-level). This two-level architecture provides the theoretical foundation for understanding how cognitive systems achieve self-regulation and control during learning activities.
According to their model, the meta-level maintains a dynamic internal representation of the object-level, functioning as a mental simulation that enables the system to monitor current cognitive states and guide transitions towards desired goals. The interaction between levels operates through two distinct information flows, control (meta-level $\rightarrow$ object-level) and monitoring (object-level $\rightarrow$ meta-level). Control processes enable the meta-level to modify object-level states or processes – such as allocating study time to difficult material or switching from rote memorisation to elaborative rehearsal strategies. Monitoring processes provide the meta-level with information about current object-level states, updating its internal model of the cognitive situation. These relationships connotes two notable properties: they are logically independent (control does not inherently generate feedback about its effects) and asymmetric (the meta-level maintains a model of the object-level whilst the object-level operates without any corresponding representation of the meta-level).
B.1 Acquisition Process
Initialisation Given a task ($\mathcal{T}$) and goal ($\mathcal{G}$) with a target performance level ($\rho^{*}$), the agent establishes the norm of study ($\mathcal{N}_{s}$):
$\mathcal{MK}_{0}^{\text{STM}}$ $=\text{retrieve}_{\theta}(\mathcal{MK},\mathcal{T},\mathcal{G})$
$\delta_{\text{retention}}$ $=\text{formulate}(\mathcal{MK}_{0}^{\text{STM}},\tau_{\text{delay}},\mathcal{T},\mathcal{G})$
$\mathcal{N}_{s}$ $=\rho^{*}\times(1+\delta_{\text{retention}})$
At the initialisation stage ($\tau=0$), a global metacognitive parameter ($\mathcal{N}_{s}$) operationalises abstract goals into quantified mastery criteria, which Nelson and Narens [1990, p. 130] define as ‘the overall degree of mastery the person believes should be attained during acquisition’.
Following Ericsson and Simon [1984], monitoring operations occur within working memory (STM), with $\mathcal{MK}_{0}^{\text{STM}}$ denoting the metacognitive knowledge retrieved into this workspace at $\tau=0$. Information from long-term memory (LTM) may be accessed by first copying it into STM with probability $\theta$ [Atkinson and Shiffrin, 1968], captured through the notation $\text{retrieve}_{\theta}(\cdot)$ for this probabilistic access during metacognitive monitoring. The term $\delta_{\text{retention}}$ represents the agent’s theory of retention – beliefs about memory decay over the interval $\tau_{\text{delay}}$.
This formulation reflects Nelson and Narens’ insight that effective learning requires anticipatory compensation for memory decay. The model predicts systematic variation in norm-setting behaviour across agents and contexts. For instance, an agent targeting 90% test performance ($\rho^{*}=0.9$) who expects 20% decay ($\delta_{\text{retention}}=0.2$) must achieve 108% mastery during acquisition. Moreover, the framework anticipates differential standards across learning contexts: conceptual understanding tasks ($G_{1}$, with $\delta_{\text{retention}}=0.1$) vs. verbatim recall tasks ($G_{2}$, with $\delta_{\text{retention}}=0.2$) yield distinct acquisition targets (99% vs. 108% respectively) even under identical performance goals.
M-G-V (Learning) Cycle For learning cycles $\tau\in\{1,\ldots,T_{\text{learn}}\}$, let $\mathcal{T}_{\tau}=\{i_{j}:j\in\mathcal{J}_{\tau}\}$ denote the set of items remaining in the task at cycle $\tau$, where $\mathcal{J}_{\tau}\subseteq\{1,2,\ldots,N\}$ represents the indices of items still requiring learning. $\Phi_{\tau}^{\text{STM}}$ represent the cumulative learning experience in working memory.
WHILE $\mathcal{S}_{\tau}=\text{ACTIVE}$:
  1. 1.
    MONITOR: Assess current mastery for each item $i_{j}\in\mathcal{T}_{\tau}$.
    Monitoring involves retrieving metacognitive knowledge and generating metacognitive experiences about the current learning state.
    $\mathcal{MK}_{\tau}^{\text{STM}}$ $=\mathcal{MK}_{\tau-1}^{\text{STM}}\cup\text{retrieve}_{\theta}(\mathcal{MK},\mathcal{ME}_{\tau-1})\quad\text{if }\tau>0$
    $\mathcal{ME}_{\tau,j}$ $=\begin{cases}[\text{EOL}_{\tau,j},\text{null}]&\text{if }\tau=0\\ [\text{FOK}_{\tau,j},\text{null}]&\text{if }\tau>0\end{cases}$
    where:
    $\text{EOL}_{\tau,j}$ $=\text{feel}(i_{j})\oplus\text{assess}(i_{j},\mathcal{MK}_{\tau}^{\text{STM}})$ $\text{if }\tau=0$
    $\text{FOK}_{\tau,j}$ $=\text{feel}(i_{j},\mathcal{CO}_{\tau-1,j})\oplus\text{assess}(i_{j},\mathcal{CO}_{\tau-1,j},\mathcal{MK}_{\tau}^{\text{STM}})$ $\text{if }\tau>0$
    Metacognitive experiences are represented as vectors, reflecting Nelson and Narens’ proposal of their multidimensional nature. Both Ease of Learning (EOL) and Feeling of Knowing (FOK) are immediate phenomenological experiences that emerge during cognitive tasks, illustrating how subjective feelings support monitoring functions, serving as the primary input for subsequent control decisions [Nelson and Narens, 1990, p. 160].
  2. 2.
    GENERATE: Transforms monitoring outputs into executable learning actions.
    Resources are allocated inversely proportional to their EOL or FOK, and strategy selection integrates metacognitive inputs to map learning methods to individual items.
    $r_{\tau,j}$ $=R_{\text{total}}\times\frac{w_{j}}{\sum_{k=1}^{N}w_{k}},\quad\text{where }w_{j}=(\mathcal{ME}_{\tau,j}[1])^{-1}$
    $\sigma_{\tau,j}$ $=\text{select}(s\in\mathcal{MK}_{\text{Strategy}}\mid i_{j},r_{\tau,j},\mathcal{ME}_{\tau,j},\mathcal{MK}_{\tau})$
    The learning plan $\mathcal{P}_{\tau,j}=(i_{j},r_{\tau,j},\sigma_{\tau,j})$ is executed to produce cognitive outcomes (new memory state).
    $\mathcal{CO}_{\tau,j}=\text{execute}(\mathcal{P}_{\tau,j})$
  3. 3.
    VERIFY: Assess learning progress and determines cycle continuation.
    Judgements of Learning (JOL) evaluate current mastery levels following cognitive outcomes.
    $\text{JOL}_{\tau,j}$ $=\text{feel}(i_{j},\mathcal{CO}_{\tau,j})\oplus\text{assess}(i_{j},\mathcal{CO}_{\tau,j},\mathcal{MK}_{\tau})$
    $\mathcal{ME}_{\tau,j}$ $=[\mathcal{ME}_{\tau,j}[1],\text{JOL}_{\tau,j}]$
    $\Phi_{\tau,j}^{\text{STM}}$ $=(\mathcal{ME}_{\tau,j},i_{j},r_{\tau,j},\sigma_{\tau,j},\mathcal{CO}_{\tau,j})$
    $\Phi_{\tau}^{\text{STM}}$ $=\Phi_{\tau-1}^{\text{STM}}\cup\{\Phi_{\tau,j}^{\text{STM}}:j\in\mathcal{J}_{\tau}\}$
    For each item $i_{j}\in\mathcal{T}_{\tau}$, an agent computes the mastery discrepancy. Items that have reached the norm of study are removed from further consideration. Thus, $\mathcal{T}_{\tau+1}=\{i_{j}:j\in\mathcal{J}_{\tau+1}\}$ contains only items still requiring learning.
    $\Delta_{\tau,j}$ $=\mathcal{N}_{s}-\text{JOL}_{\tau,j}$
    $\mathcal{J}_{\tau+1}$ $=\{j\in\mathcal{J}_{\tau}:\Delta_{\tau,j}>0\}$
    Learning continues as long as any item remains below the mastery threshold:
    $\mathcal{S}_{\tau+1}=\begin{cases}\text{ACTIVE}&\text{if }\mathcal{J}_{\tau+1}\neq\emptyset\\ \text{TERMINATE}&\text{otherwise}\end{cases}$
B.2 Retrieval Process
Initialisation Given a retrieval query $\mathcal{Q}$, the agent establishes retrieval goals and accesses contextually relevant metacognitive knowledge for search control.
$\mathcal{MK}_{0}^{\text{STM}}$ $=\text{retrieve}_{\theta}(\mathcal{MK},\mathcal{Q})$
Nelson and Narens [1990] conceptualise FOK through the dual-counter hypothesis: one component accumulates evidence for information presence in memory (affirmative FOK, $\text{FOK}_{\tau}^{+}$), while the other accumulates evidence for information absence, consistent with ‘knowing not’ [Kolers and Palef, 1976] (negative FOK, $\text{FOK}_{\tau}^{-}$). This dual mechanism enables both continued search when positive evidence accumulates and efficient termination when negative evidence dominates, preventing exhaustive search behaviour.
The initial thresholds $\lambda_{\text{confidence}}^{(0)}$ and $\lambda_{\text{FOK}}^{(0)}$ are established through the agent’s privileged access to personal metacognitive calibration history within $\mathcal{MK}_{0}^{\text{STM}}$:
$\lambda_{\text{FOK}}^{(0)}$ $=\text{median}(\{||\text{FOK}||:\text{successful retrievals in }\mathcal{MK}_{0}^{\text{STM}}\})$
[…]

### v3

@@ 3.2.3 Memory Consolidation and Knowledge Evolution
4 Limitations and Future Work
4.1 From Specification to Implementation
The MGV framework specifies what should be computed but not how to compute it in neural architectures. Operationalising constructs like $\mathcal{ME}_{\text{difficulty}}$ or dual-counter FOK requires identifying appropriate neural correlates or designing explicit computational mechanisms. For language models, this might involve obtaining logit-based confidence ratings (Kumaran et al., 2025), using entropy-based proxies for metacognitive experiences, or implementing explicit evidence accumulators. Recent work suggests that LLMs access metacognitive signals they cannot adequately articulate, with implicit confidence measures outperforming explicit verbalisation (Wang et al., 2025; Tian et al., 2023; Xiong et al., 2023; Cash et al., 2024; Griot et al., 2025; Lindsey et al., 2025). Neural analysis reveals a lower-dimensional “metacognitive space” where monitoring signals correspond to linearly separable directions (Ji-An et al., 2025; Zou et al., 2023; Liu et al., 2023), consistent with cognitive findings that metacognition operates on abstracted representations (Reder, 1987).
4.2 Scope of Formalisation
The present work formalises two foundational theories, but metacognition research extends well beyond Flavell and Nelson and Narens. A more complete computational account of metacognition should integrate subsequent theoretical developments. Additionally, our formalisations make interpretive choices where the source theories are ambiguous or silent. For instance, the exponential decay of satisficing thresholds reflects our reading of the theories rather than explicit specifications in the original texts. Alternative formalisations are possible and might yield different architectural implications.
4.3 Normative Grounding
The framework lacks principled justification for its architectural choices. Resource-rational analysis explains why certain metacognitive policies are optimal given computational constraints. MGV, by contrast, specifies mechanisms without explaining why those mechanisms are appropriate. As Callaway et al. (2024, pp. 17–18) noted regarding related work in metamemory, this paper shares the limits of Metcalfe (1993), Koriat and Goldsmith (1996), and Bennett et al. (2017) in focusing on functional aspects without taking an explicitly optimal approach. That is, while resource-rational analysis could potentially explain MGV-like mechanisms as approximations to optimal behaviour, MGV cannot explain resource-rational optima as emergent from its architecture.
Yet this limitation is a consequence of methodological choice rather than oversight. We prioritised preserving psychological detail over deriving mechanisms from first principles. The result is a framework that describes what metacognitive components might look like but cannot explain why they should take those forms rather than others. Whether dual-counter FOK approximates optimal stopping behaviour, whether satisficing thresholds should decay exponentially rather than linearly, whether tripartite knowledge organisation is functionally necessary: these questions cannot be answered within MGV but could potentially be addressed through resource-rational analysis. The grounding of MGV in resource-rational principles represents one path toward resolving this limitation.
4.4 Future Directions
Grounding MGV in Resource-Rational Principles
The most productive path forward may involve selective integration of the two approaches. Some MGV constructs appear directly amenable to resource-rational analysis. The stopping dynamics formalised in Nelson and Narens’ retrieval model parallel the optimal stopping structure of memory recall in the resource-rational framework, where feeling-of-knowing serves as the monitoring signal that enables efficient termination of searches unlikely to succeed (Callaway et al., 2024). Flavell’s cognitive monitoring, which involves inferring the state of ongoing cognitive processes from imperfect signals, shares the partial observability structure of the belief-MDP formulation of cognitive control (Lieder et al., 2018). The threshold dynamics that govern when to cease retrieval or when to shift strategies may be derivable from VOC calculations rather than stipulated as free parameters.
Other constructs may serve primarily descriptive functions that complement rather than compete with normative analysis. The tripartite knowledge taxonomy (person, task, and strategy knowledge) provides a vocabulary for the features over which resource-rational agents must learn, but the taxonomy itself is not derived from optimality principles. Similarly, the phenomenological vocabulary of metacognitive experiences (feeling-of-knowing, judgment-of-learning, feeling-of-confidence) names the signals that enable metacognitive control, but resource-rational analysis explains why such signals are useful rather than what they feel like. A complete theory may require both the normative grounding that resource-rational analysis provides and the psychological granularity that MGV preserves.
Meta-Reasoning and Meta-Learning in Language Models.
[…]